The optimal solution for the following LPP is Z* = 30 :

Max Z = X1 - X2 + 3X3

Subject to X1 + X2 + X3 less than or equal to 10

X1, X2, X3 greater than or equal to zero .


1
Expert's answer
2021-05-11T18:26:16-0400

Given, Max "Z=x_1-x_2+3x_3"


subject to- "x_1+x_2+x_3\le 10~~~~~-(1)"

"x_1,x_2,x_3\ge 0"


Z=30( given)

"\Rightarrow x_1-x_2+2x_3=30~~~~~~~~~-(2)"


The solution to the above equation equation can be find by calculating optimal solution.


Solving eqn.(1) and (2) by taking "x_3" as restricted variable , SO Taking "x_3" =0


we get "x_1=20, x_2=-10,x_3=0" .